OpenStudy (anonymous):

Okay, I've found my greatest weakness... SLOPES! Can anyone help me figure out how to do this problem? ]Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

x intercept = (8, 0); y intercept = (0, -4)

OpenStudy (anonymous):

two methods Ax + By = 1 or y = mx +b

OpenStudy (anonymous):

Dankeschön! I need to use y=mx+b! :)

OpenStudy (anonymous):

i like y = mx + b I solve for y then b is the y intercept then I input 0 for the y value and solve for x y = 1/2 x - 4 0 = 1/2 x - 4 therefore x = 8

OpenStudy (anonymous):

the other way, Ax + By = 1 is faster if you understand it; however, I was taught y = mx +b first

OpenStudy (anonymous):

I was taught y=mx+b as the first way, and the standard version was too hard for me! Thanks!